# A Cocktail-Party Benchmark: Multi-Modal dataset and Comparative Evaluation Results

### v1

@@ 1 Introduction
In addition to research on acoustic robustness, audio-visual speech recognition (AVSR) has emerged as a promising approach by leveraging visual cues such as lip movements to complement the acoustic signal [10, 11, 12, 13, 14]. To advance this direction, several audio-visual and multi-modal datasets have been introduced. Large-scale corpora such as LRS2/LRS3 [15, 16], Muavic [17] focus primarily on single-speaker or Ego4D [18] broad egocentric scenarios, while conversational datasets such as AMI [9], MISP [19], and MMCSG [20] capture multi-party meetings but are limited to single conversations per session. CHiME-5/6 [5] capture multi-party meetings but remain limited to audio-only. Despite these efforts, there remains a clear gap, audio-visual, natural, unscripted, multi-party interaction under cocktail-party conditions. The challenge in such settings arises from the so-called cocktail-party phenomenon, which describes humans’ ability to selectively attend to one conversation while ignoring others. In this context, systems must determine not only who speaks when and what is spoken, but also with whom the interaction occurs.
The MCoRec task aims to bridge the gap between previous efforts by establishing a benchmark for the cocktail-party challenge. Our dataset features audio and video recordings of up to eight participants engaged in up to four simultaneous conversations. The challenge provides a single 360° video and a single-channel audio recording, reflecting typical conditions using commercially available devices. Benchmarking systems will focus on jointly transcribing speech and clustering speakers into conversation groups. The MCoRec dataset and baseline systems have been made reproducible via https://github.com/MCoRec/mcorec_baseline
@@ 4.1 Active Speaker Detection (ASD)
On the development set, the model achieves a Micro-Averaged Intersection over Union (IoU) of 75.58%. IoU simply measures how much the predicted active segments overlap with the annotated ground-truth segments. Although this metric is not used for benchmarking systems in the challenge, it provides useful insight into the performance of the ASD module as an intermediate step. A value of 75.58% indicates that the model can localize active speaking regions with reasonably accuracy, but also leaves room for improvement, especially in challenging multi-speaker scenarios.

### v2

@@ 1 Introduction
In addition to research on acoustic robustness, audio-visual speech recognition (AVSR) has emerged as a promising approach by leveraging visual cues such as lip movements to complement the acoustic signal [12, 21, 20, 5, 26, 16, 11]. To advance this direction, several audio-visual and multi-modal datasets have been introduced. Large-scale corpora such as LRS2/LRS3 [9, 1] and Muavic [3] focus primarily on single-speaker or Ego4D [13] broad egocentric scenarios, while conversational datasets such as AMI [7], MISP [8], and MMCSG [28] capture multi-party meetings but are limited to single conversations per session. CHiME-5/6 [25] capture multi-party meetings but remain limited to audio-only. Despite these efforts, there remains a clear gap: audio-visual, natural, unscripted, multi-party interaction under cocktail-party conditions. The challenge in such settings arises from the so-called cocktail-party phenomenon, which describes humans’ ability to selectively attend to one conversation while ignoring others. In this context, systems must determine not only who speaks when and what is spoken, but also with whom the interaction occurs.
The MCoRec task aims to bridge the gap between previous efforts by establishing a benchmark for the cocktail-party challenge. Our dataset features audio and video recordings of up to eight participants engaged in up to four simultaneous conversations. The challenge provides a single 360° video and a single-channel audio recording, reflecting typical conditions using commercially available devices. Benchmarking systems will focus on jointly transcribing speech and clustering speakers into conversation groups.
@@ 2 Dataset
The MCoRec dataset was developed externally by Interactive-AI LLC and is made available for the CHiME Challenge for research purposes. For commercial use, please contact authors. Dataset and baseline systems can be accessed via https://github.com/MCoRec/mcorec_baseline.
@@ 4.1 Active Speaker Detection (ASD)
On the development set, the model achieves a Micro-Averaged Intersection over Union (IoU) of 75.58%. IoU measures how much the predicted active segments overlap with the annotated ground-truth segments. Although this metric is not used for benchmarking systems in the challenge, it provides useful insight into the performance of the ASD module as an intermediate step. A value of 75.58% indicates that the model can localize active speaking regions with reasonable accuracy, but also leaves room for improvement, especially in challenging multi-speaker scenarios.
@@ 5 Conclusion
6 Acknowledgment
The authors gratefully acknowledge support from the EU’s Horizon research & innovation programme (101135798 – Meetween; 101213369 – DVPS). We thank Interactive-AI LLC for providing the database for research purposes.